A random sample of 35 used car lots found that the average price of a used car is $11,750. Assume the population standard deviation is $1,225. Using a 90% confidence level, calculate the following:

Critical Value

Group of answer choices A.1.645 B. 1.96 C . 2.05 D. 1.28

* please show me how you worked it out. Thanks!
Answer #1

solution

At 90% confidence level

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645 ( Using z table )

critical value=1.645